College of Computer & Information Science Fall 2004 Northeastern University Handout 1 CSU640: Network Fundamentals 13 September 2004

Problem Set 1 (due September 23)

1. (10 points) Propagation and transmission delays

Problem 6, page 64.

2. (10 points) File transfer time

Calculate the total time required to transfer a 1000 KB file in the following cases, assuming an RTT of 100 ms, a packet size of 1 KB data, and an initial 2 × RTT of “handshaking” before data is sent. Assume that the RTT time includes both propagation and queuing delays. The file transmission is complete when the entire file is received at the receiver.

(a) The bandwidth is 1.5 Mbps, and data packets are sent continuously.

(b) The bandwidth is 1.5 Mbps, but we send only one packet per RTT; that is, after we finish sending each data packet we must wait one RTT before sending the next.

(c) The bandwidth is infinite, meaning that we take transmit time to be zero, and up to 20 packets can be sent per RTT.

(d) The bandwidth is infinite, and during the first RTT we can send one packet (2^{0}), during the second RTT we can send two packets (2^{1}), during the third we can send four packets (2^{2}), and so on. That is, in the ith RTT, we can send 2^{i−1} packets.

3. (10 points) Message segmentation

Problem 20, pages 66-67.

4. (10 points) Traceroute

Problem 13, page 65. Attach outputs of your traceroutes and the times they were executed. For generating the routes, you may use the website www.traceroute.org, or the traceroute command: traceroute in Unix/Linux, tracert in Windows.